IF f(4)=0 and f(6)=6 find f(x).
Answer by rajagopalan(148)

dear student
given f(4)=0.....c=0
let the fn be ax^2+bx+c
f(4)=16a+4b=0...eqn 1
f(6)=36a+6b=6...eqn 2...6a+b=1
solving
16a+4b=0
6a+b=1
16a+4b=0
24a+4b=4 (eqn 2*4)
subtract 8a=4
a=1/2
b=1-6a=1-3=-2
so the requsite functin = (x^2/2)-2x
verification
f(4)=4^2/2 -2*4 = 8-8=0
and
f(6)=6^2/2 - 6*2 =18-12=6
Okay